OpenStudy (henrietepurina):

Algebra 1 question :D Medal + Fan!

OpenStudy (henrietepurina):

A ball is thrown upward from the top of a building. The function below shows the height of the ball in relation to sea level, ft, in feet, at different times t in seconds: f(t) = -16t^2 + 48t + 100 The average rate of change of f(t) from t = 3 seconds to t = 5 seconds is _____feet per second.

OpenStudy (misty1212):

HI!!!

OpenStudy (henrietepurina):

heya :D

OpenStudy (misty1212):

you never quit doing math!

OpenStudy (swissgirl):

#MathisLyfe

OpenStudy (henrietepurina):

no, I work on it all day :)

OpenStudy (henrietepurina):

really! :D

OpenStudy (misty1212):

for this problem we need a couple numbers we are going to compute \[\frac{f(5)-f(3)}{5-3}\]

OpenStudy (henrietepurina):

that's how I spend my weekends practically :D

OpenStudy (henrietepurina):

ok, i'll try doing that :)

OpenStudy (misty1212):

so we need to find \(f(5)\) and \(f(3)\) because we already know that \(5-3=2\)

OpenStudy (henrietepurina):

f(2)/2

OpenStudy (henrietepurina):

f is the answer?

OpenStudy (misty1212):

oh no dear you could not be farther from the truth first off, \(f(5)f(3)\neq f(2)\) we have to compute those numbers

OpenStudy (henrietepurina):

ah ok then :D

OpenStudy (misty1212):

secondly, \(f\) is a function, not a variable, so \(\frac{f(2)}{2}\neq f\) that is just silly

OpenStudy (henrietepurina):

ok sorry :|

OpenStudy (misty1212):

you have \(f(t) = -16t^2 + 48t + 100\) so what is \(f(5)\)? feel free to use a calculator

OpenStudy (henrietepurina):

give me a second ok?

OpenStudy (misty1212):

take your time

OpenStudy (henrietepurina):

umm... could it be −6060?

OpenStudy (misty1212):

idk let me check

OpenStudy (henrietepurina):

kk

OpenStudy (misty1212):

that is not even close to what i got \[f(t) = -16t^2 + 48t + 100\] \[f(5)=-16\times 5^2+48\times 5+100\] use a calculator

OpenStudy (henrietepurina):

I did... ohhh! 16*((5)^2) not ((16*5)^2)

OpenStudy (misty1212):

or better yet, use this http://www.wolframalpha.com/input/?i=+-16*5^2+%2B+48*5+%2B+100

OpenStudy (henrietepurina):

-60 it is

OpenStudy (misty1212):

lol yeah don't forget the old "order o' operations"

OpenStudy (henrietepurina):

sorry, messed up on my behalf

OpenStudy (misty1212):

that's ok dear, let me know when you get \(-60\)

OpenStudy (henrietepurina):

yeah, I already typed that I got -60

OpenStudy (misty1212):

ok now then next number we need is \(f(2)\) think we can get that ?

OpenStudy (henrietepurina):

yes I can

OpenStudy (henrietepurina):

132

OpenStudy (misty1212):

got it !!

OpenStudy (henrietepurina):

yay...

OpenStudy (misty1212):

TOO BAD I MADE A MISTAKE

OpenStudy (henrietepurina):

what do you mean?

OpenStudy (misty1212):

lol we need \(f(3)\) not \(f(2)\) doe

OpenStudy (henrietepurina):

it's ok i'll redo it :D

OpenStudy (henrietepurina):

100

OpenStudy (misty1212):

we got the \(f(5)\) so that is ok, we need \(f(3)\) let me check

OpenStudy (henrietepurina):

its 100

OpenStudy (misty1212):

\[100\] it is !

OpenStudy (henrietepurina):

ok then :)

OpenStudy (misty1212):

final job is \[\frac{-60-100}{2}\]

OpenStudy (henrietepurina):

haha ok...

OpenStudy (misty1212):

i.e. \[\frac{f(5)-f(3)}{5-3}\]

OpenStudy (henrietepurina):

i got -80

OpenStudy (misty1212):

once we compute it and get the correct answer, we have to have a brief explanation of why this question was written by a moron who knows no math and should not be writing questions for unsuspecting students

OpenStudy (misty1212):

ok good, \(-80\) is correct at least that is what the idiot wants do you have an actual teacher for this class, or is it on line

OpenStudy (henrietepurina):

class

OpenStudy (misty1212):

so you have a real teacher?

OpenStudy (henrietepurina):

yeah

OpenStudy (misty1212):

ok explain to this teacher (who may be a very nice person, but should be teaching gym) that if the question really is A ball is thrown upward from the top of a building. The function below shows the height of the ball in relation to sea level, ft, in feet, at different times t in seconds: then the ball hits the ground before 5 seconds has elapsed, ( we know this since \(f(5)=-60\) ) so it cannot have been in the air for five seconds

OpenStudy (henrietepurina):

(don't worry, she's not one of the best)

OpenStudy (misty1212):

in other words it makes no sense to say the ball is \(-60\) when it hits the ground, \(h=0\) then presumably it does not go underground

OpenStudy (henrietepurina):

mhm... so would it be correct to say -60 anyway?

OpenStudy (misty1212):

go ahead and ask how the ball can be \(-60\) feet see what she says

OpenStudy (henrietepurina):

wait wait, when I divided again, I got -80, but that doesn't make a difference

OpenStudy (misty1212):

believe me the answer she is looking for is \[\frac{f(5)-f(3)}{5-3}=\frac{-60-100}{2}=-80\]

OpenStudy (misty1212):

all i am saying is that the ball does not stay in the air for five seconds, so the word problem aspect of this is stupid if it was just "find the average rate of change of \(f(t)=-16t^2+48t+100\) then \(-80\) makes perfect sense

OpenStudy (misty1212):

this is what happens when not too sophisticated math teacher try to make up word problems i would be better if they just stuck to equations and such

OpenStudy (henrietepurina):

thanks so much... sorry for not responding, BATHROOM BREAK!

OpenStudy (henrietepurina):

wait wait wait @misty1212 .... remember that this question is asking for the rate of change, so wouldn't the answer actually be logical?

OpenStudy (misty1212):

no, not as a word problem as a math problem, yes, the average rate of change of that function on the interval \((3,5)\) is \(-80\)

OpenStudy (henrietepurina):

ah yeah, I see where you are going :D

OpenStudy (misty1212):

as a word problem it does not make sense because the ball cannot be 60 feet under ground

OpenStudy (henrietepurina):

yeah yeah :D BTW... mind helping me on another one?

OpenStudy (misty1212):

as a pure math problem forgetting the words, then yes the rate of change is -80

OpenStudy (misty1212):

sure why not but post in a different thread it gets hard to scroll down so much

OpenStudy (henrietepurina):

ok can do :D

OpenStudy (anonymous):

was this correct?